OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2

OpenStudy (anonymous):

=4(-2) - (3) -2(-2) =-8-3+4 =-4-3 =-7

OpenStudy (ryan123345):

\[4(-2) - 3 - 2(-2)\]

OpenStudy (ryan123345):

all you have to do is plug in the numbers and solve. Make sure you use PEMDAS properly.